### v1

@@ Appendix E Limitations and Future Work
Qwen2-VL-72B-Instruct 0.572 0.528 0.502 0.525 0.578 0.559 0.526 0.562 0.563 0.549 0.567 0.603 0.548
qwen-vl-max-0809 0.568 0.520 0.506 0.536 0.573 0.558 0.521 0.554 0.553 0.541 0.572 0.608 0.538
InternVL2-76B 0.616 0.572 0.540 0.597 0.614 0.611 0.569 0.604 0.594 0.581 0.608 0.640 0.588
GPT4o_HIGH 0.581 0.543 0.516 0.546 0.576 0.574 0.533 0.581 0.567 0.558 0.586 0.620 0.568
InternVL2-40B 0.584 0.527 0.518 0.528 0.592 0.577 0.528 0.581 0.563 0.560 0.574 0.607 0.565
Ovis1.6-Gemma2-9B 0.601 0.547 0.518 0.573 0.593 0.593 0.540 0.595 0.579 0.568 0.580 0.633 0.576
llava_onevision_qwen2_72b_ov 0.512 0.463 0.441 0.463 0.518 0.500 0.458 0.504 0.494 0.478 0.510 0.544 0.509
llava_onevision_qwen2_72b_si 0.530 0.477 0.457 0.500 0.540 0.518 0.467 0.512 0.517 0.515 0.527 0.559 0.509
Qwen2-VL-7B-Instruct 0.550 0.492 0.480 0.514 0.556 0.539 0.488 0.540 0.535 0.515 0.542 0.584 0.520
InternVL2-26B 0.596 0.547 0.522 0.562 0.586 0.583 0.538 0.587 0.572 0.569 0.581 0.627 0.564
qwen-vl-plus 0.497 0.455 0.440 0.386 0.516 0.486 0.457 0.444 0.467 0.464 0.496 0.484 0.486
MiniCPM-V-2_6 0.559 0.517 0.492 0.524 0.565 0.559 0.500 0.577 0.540 0.547 0.558 0.594 0.522
InternVL2-8B 0.564 0.522 0.501 0.526 0.576 0.557 0.507 0.539 0.545 0.536 0.560 0.596 0.543
Ovis1.5-Gemma2-9B 0.541 0.493 0.470 0.490 0.546 0.526 0.483 0.541 0.535 0.515 0.532 0.557 0.528
RBDash_72b 0.533 0.499 0.491 0.497 0.545 0.531 0.506 0.549 0.540 0.514 0.533 0.553 0.512
MMAlaya2 0.557 0.511 0.485 0.510 0.554 0.549 0.497 0.554 0.524 0.534 0.549 0.583 0.531
Ovis1.5-Llama3-8B 0.540 0.505 0.485 0.499 0.543 0.526 0.489 0.530 0.539 0.518 0.540 0.565 0.520
OmChat 0.459 0.415 0.394 0.365 0.455 0.437 0.432 0.450 0.457 0.449 0.449 0.468 0.432
InternVL-Chat-V1-5 0.569 0.522 0.493 0.518 0.562 0.556 0.519 0.556 0.534 0.546 0.559 0.581 0.545
llava_onevision_qwen2_7b_si 0.514 0.462 0.439 0.454 0.520 0.498 0.446 0.494 0.504 0.485 0.514 0.543 0.485
XComposer2d5 0.446 0.449 0.470 0.411 0.472 0.458 0.472 0.410 0.459 0.441 0.482 0.472 0.478
Pixtral-12B 0.567 0.525 0.513 0.530 0.574 0.557 0.518 0.557 0.554 0.550 0.561 0.579 0.547
InternVL2-4B 0.573 0.524 0.507 0.534 0.578 0.564 0.525 0.540 0.546 0.540 0.574 0.587 0.552
llava_onevision_qwen2_7b_ov 0.504 0.452 0.439 0.470 0.512 0.496 0.432 0.491 0.491 0.465 0.513 0.541 0.478
glm-4v-9b 0.534 0.487 0.474 0.413 0.544 0.522 0.476 0.490 0.490 0.504 0.546 0.536 0.519
molmo-7B-D-0924 0.537 0.495 0.483 0.447 0.550 0.529 0.496 0.509 0.524 0.507 0.535 0.551 0.518
XComposer2_4KHD 0.444 0.419 0.388 0.387 0.457 0.433 0.425 0.417 0.428 0.435 0.465 0.469 0.437
MiniCPM-Llama3-V-2_5 0.539 0.498 0.478 0.485 0.547 0.534 0.496 0.523 0.529 0.526 0.539 0.555 0.521
VILA1.5-40b 0.500 0.443 0.417 0.448 0.495 0.492 0.447 0.508 0.458 0.482 0.506 0.536 0.497
cambrian_34b 0.567 0.516 0.503 0.517 0.567 0.552 0.508 0.553 0.535 0.548 0.554 0.592 0.539
WeMM 0.508 0.465 0.445 0.437 0.526 0.493 0.447 0.496 0.495 0.467 0.510 0.519 0.477
Llama-3.2-11B-Vision-Instruct 0.541 0.510 0.501 0.491 0.544 0.529 0.513 0.520 0.513 0.533 0.543 0.561 0.530
Qwen2-VL-2B-Instruct 0.505 0.462 0.439 0.456 0.519 0.489 0.456 0.486 0.495 0.475 0.494 0.539 0.486
XComposer2 0.264 0.244 0.236 0.175 0.268 0.257 0.207 0.243 0.276 0.255 0.265 0.279 0.249
cogvlm2-llama3-chat-19B 0.486 0.455 0.445 0.382 0.506 0.467 0.440 0.437 0.458 0.461 0.496 0.490 0.464
Idefics3-8B-Llama3 0.479 0.428 0.403 0.414 0.479 0.465 0.437 0.486 0.447 0.454 0.484 0.511 0.448
Mini-InternVL-Chat-4B-V1-5 0.541 0.496 0.484 0.488 0.551 0.533 0.489 0.534 0.529 0.522 0.551 0.562 0.525
XinYuan-VL-2B-Instruct 0.393 0.378 0.365 0.337 0.427 0.392 0.387 0.379 0.395 0.393 0.416 0.414 0.398
llava_next_yi_34b 0.498 0.451 0.432 0.391 0.511 0.485 0.429 0.480 0.475 0.472 0.489 0.515 0.461
InternVL2-2B 0.501 0.450 0.440 0.395 0.511 0.479 0.440 0.469 0.479 0.453 0.501 0.498 0.471
Phi-3-Vision 0.469 0.431 0.423 0.365 0.493 0.464 0.442 0.418 0.457 0.449 0.489 0.485 0.473
cambrian_13b 0.535 0.474 0.468 0.465 0.540 0.508 0.477 0.502 0.526 0.521 0.532 0.546 0.519
Phi-3.5-Vision 0.499 0.447 0.428 0.448 0.502 0.485 0.429 0.478 0.474 0.479 0.498 0.524 0.478
idefics2_8b 0.467 0.425 0.408 0.402 0.488 0.451 0.402 0.431 0.443 0.459 0.474 0.486 0.451
cambrian_8b 0.529 0.485 0.460 0.458 0.537 0.520 0.463 0.518 0.518 0.504 0.533 0.541 0.512
minimonkey 0.467 0.422 0.413 0.374 0.488 0.455 0.441 0.430 0.459 0.429 0.478 0.465 0.454
Llama-3-MixSenseV1_1 0.492 0.448 0.423 0.420 0.489 0.484 0.434 0.476 0.481 0.466 0.490 0.526 0.459
llava_next_qwen_32b 0.526 0.479 0.456 0.449 0.539 0.516 0.468 0.494 0.514 0.489 0.528 0.543 0.501
xgen-mm-phi3-interleave-r-v1.5 0.549 0.503 0.492 0.495 0.550 0.533 0.515 0.554 0.529 0.511 0.549 0.562 0.524
Eagle-X5-13B 0.446 0.420 0.390 0.387 0.455 0.440 0.407 0.433 0.424 0.440 0.458 0.443 0.439
xgen-mm-phi3-dpo-r-v1.5 0.420 0.399 0.391 0.289 0.464 0.414 0.402 0.409 0.412 0.430 0.468 0.438 0.415
Mini-InternVL-Chat-2B-V1-5 0.447 0.408 0.394 0.308 0.474 0.433 0.412 0.410 0.430 0.416 0.455 0.439 0.426
llava_next_llama3 0.495 0.455 0.427 0.426 0.514 0.485 0.441 0.473 0.478 0.474 0.493 0.516 0.474
VILA1.5-13b 0.499 0.458 0.432 0.473 0.496 0.491 0.468 0.517 0.468 0.484 0.502 0.508 0.486
Mantis-8B-siglip-llama3 0.424 0.408 0.376 0.365 0.418 0.409 0.411 0.406 0.401 0.433 0.441 0.421 0.422

### v2

@@ Appendix E Limitations and Future Work
Qwen2-VL-72B-Instruct [65] 0.572 0.528 0.502 0.525 0.578 0.559 0.526 0.562 0.563 0.549 0.567 0.603 0.548
qwen-vl-max-0809 [6] 0.568 0.520 0.506 0.536 0.573 0.558 0.521 0.554 0.553 0.541 0.572 0.608 0.538
GPT4o_HIGH [53] 0.581 0.543 0.516 0.546 0.576 0.574 0.533 0.581 0.567 0.558 0.586 0.620 0.568
llava_onevision_qwen2_72b_ov [37] 0.512 0.463 0.441 0.463 0.518 0.500 0.458 0.504 0.494 0.478 0.510 0.544 0.509
llava_onevision_qwen2_72b_si [37] 0.530 0.477 0.457 0.500 0.540 0.518 0.467 0.512 0.517 0.515 0.527 0.559 0.509
Qwen2-VL-7B-Instruct [65] 0.550 0.492 0.480 0.514 0.556 0.539 0.488 0.540 0.535 0.515 0.542 0.584 0.520
MiniCPM-V-2_6 [75] 0.559 0.517 0.492 0.524 0.565 0.559 0.500 0.577 0.540 0.547 0.558 0.594 0.522
RBDash_72b [57] 0.533 0.499 0.491 0.497 0.545 0.531 0.506 0.549 0.540 0.514 0.533 0.553 0.512
MMAlaya2 [48] 0.557 0.511 0.485 0.510 0.554 0.549 0.497 0.554 0.524 0.534 0.549 0.583 0.531
OmChat [82] 0.459 0.415 0.394 0.365 0.455 0.437 0.432 0.450 0.457 0.449 0.449 0.468 0.432
InternVL-Chat-V1-5 [11] 0.569 0.522 0.493 0.518 0.562 0.556 0.519 0.556 0.534 0.546 0.559 0.581 0.545
llava_onevision_qwen2_7b_si [37] 0.514 0.462 0.439 0.454 0.520 0.498 0.446 0.494 0.504 0.485 0.514 0.543 0.485
XComposer2d5 [11] 0.446 0.449 0.470 0.411 0.472 0.458 0.472 0.410 0.459 0.441 0.482 0.472 0.478
Pixtral-12B [3] 0.567 0.525 0.513 0.530 0.574 0.557 0.518 0.557 0.554 0.550 0.561 0.579 0.547
llava_onevision_qwen2_7b_ov [37] 0.504 0.452 0.439 0.470 0.512 0.496 0.432 0.491 0.491 0.465 0.513 0.541 0.478
glm-4v-9b [22] 0.534 0.487 0.474 0.413 0.544 0.522 0.476 0.490 0.490 0.504 0.546 0.536 0.519
molmo-7B-D-0924 [14] 0.537 0.495 0.483 0.447 0.550 0.529 0.496 0.509 0.524 0.507 0.535 0.551 0.518
XComposer2_4KHD [11] 0.444 0.419 0.388 0.387 0.457 0.433 0.425 0.417 0.428 0.435 0.465 0.469 0.437
MiniCPM-Llama3-V-2_5 [75] 0.539 0.498 0.478 0.485 0.547 0.534 0.496 0.523 0.529 0.526 0.539 0.555 0.521
VILA1.5-40b [40] 0.500 0.443 0.417 0.448 0.495 0.492 0.447 0.508 0.458 0.482 0.506 0.536 0.497
cambrian_34b [63] 0.567 0.516 0.503 0.517 0.567 0.552 0.508 0.553 0.535 0.548 0.554 0.592 0.539
WeMM [67] 0.508 0.465 0.445 0.437 0.526 0.493 0.447 0.496 0.495 0.467 0.510 0.519 0.477
Llama-3.2-11B-Vision-Instruct [52] 0.541 0.510 0.501 0.491 0.544 0.529 0.513 0.520 0.513 0.533 0.543 0.561 0.530
Qwen2-VL-2B-Instruct [65] 0.505 0.462 0.439 0.456 0.519 0.489 0.456 0.486 0.495 0.475 0.494 0.539 0.486
XComposer2 [10] 0.264 0.244 0.236 0.175 0.268 0.257 0.207 0.243 0.276 0.255 0.265 0.279 0.249
cogvlm2-llama3-chat-19B [26] 0.486 0.455 0.445 0.382 0.506 0.467 0.440 0.437 0.458 0.461 0.496 0.490 0.464
Idefics3-8B-Llama3 [35] 0.479 0.428 0.403 0.414 0.479 0.465 0.437 0.486 0.447 0.454 0.484 0.511 0.448
Mini-InternVL-Chat-4B-V1-5 [21] 0.541 0.496 0.484 0.488 0.551 0.533 0.489 0.534 0.529 0.522 0.551 0.562 0.525
XinYuan-VL-2B-Instruct [71] 0.393 0.378 0.365 0.337 0.427 0.392 0.387 0.379 0.395 0.393 0.416 0.414 0.398
llava_next_yi_34b [43] 0.498 0.451 0.432 0.391 0.511 0.485 0.429 0.480 0.475 0.472 0.489 0.515 0.461
Phi-3-Vision [1] 0.469 0.431 0.423 0.365 0.493 0.464 0.442 0.418 0.457 0.449 0.489 0.485 0.473
cambrian_13b [63] 0.535 0.474 0.468 0.465 0.540 0.508 0.477 0.502 0.526 0.521 0.532 0.546 0.519
Phi-3.5-Vision [1] 0.499 0.447 0.428 0.448 0.502 0.485 0.429 0.478 0.474 0.479 0.498 0.524 0.478
idefics2_8b [36] 0.467 0.425 0.408 0.402 0.488 0.451 0.402 0.431 0.443 0.459 0.474 0.486 0.451
cambrian_8b [63] 0.529 0.485 0.460 0.458 0.537 0.520 0.463 0.518 0.518 0.504 0.533 0.541 0.512
minimonkey [29] 0.467 0.422 0.413 0.374 0.488 0.455 0.441 0.430 0.459 0.429 0.478 0.465 0.454
Llama-3-MixSenseV1_1 [64] 0.492 0.448 0.423 0.420 0.489 0.484 0.434 0.476 0.481 0.466 0.490 0.526 0.459
llava_next_qwen_32b [43] 0.526 0.479 0.456 0.449 0.539 0.516 0.468 0.494 0.514 0.489 0.528 0.543 0.501
xgen-mm-phi3-interleave-r-v1.5 [73] 0.549 0.503 0.492 0.495 0.550 0.533 0.515 0.554 0.529 0.511 0.549 0.562 0.524
Eagle-X5-13B [59] 0.446 0.420 0.390 0.387 0.455 0.440 0.407 0.433 0.424 0.440 0.458 0.443 0.439
xgen-mm-phi3-dpo-r-v1.5 [73] 0.420 0.399 0.391 0.289 0.464 0.414 0.402 0.409 0.412 0.430 0.468 0.438 0.415
Mini-InternVL-Chat-2B-V1-5 [21] 0.447 0.408 0.394 0.308 0.474 0.433 0.412 0.410 0.430 0.416 0.455 0.439 0.426
llava_next_llama3 [43] 0.495 0.455 0.427 0.426 0.514 0.485 0.441 0.473 0.478 0.474 0.493 0.516 0.474
VILA1.5-13b [40] 0.499 0.458 0.432 0.473 0.496 0.491 0.468 0.517 0.468 0.484 0.502 0.508 0.486
Mantis-8B-siglip-llama3 [31] 0.424 0.408 0.376 0.365 0.418 0.409 0.411 0.406 0.401 0.433 0.441 0.421 0.422